Moist air enters a cooling coil at 27°C DBT and 60 percent relative humidity at a flow rate of 1.15 kg/s and leaves at 16°C DBT and 90 percent relative humidity. Calculate the latente load on the coil, the sensible heat ratio and the bypass factor.
Answer #1

. coel Inlet condition of moist air at cooling T, = 27°C dry bulb temperature 8, = 60% = Relative humidity Outlet condition Tria Chi-ha) ås from psychrometric chart, T₂ = 16°c (DBT & Ra=90% RH at h₂ = 42 kJ/kg Qs = 1.15 (53 - 42) ås = 12.65 12.65 kWTade = 13°c 16-13 BP = = 0.214 27 - 13 BPF 0.214 (4-14) ow= I specific Humidi (110 Lap 88/8) 1 Qs malni - ha) T2 Ti Tadp Dry100 105 110 115 120 kJ/kg/d.a) 30 29 29 ..0- - 0.8 -10.0 28 10.00 20 -5.0 8.0 -0.7 7.0 6.0 +0.6 SENSIBLE HEAT TOTAL HEAT -4.0100 105 110 115 120 kJ/kg/d.a) 30 95 29 -1.0- .- 00 8:9 10.007 -0.8 -10.0 1.5 +5.0 2.0 27 1.6 6.0 - 1.0 SENSIBLE HEAT TOTAL H